OpenStudy (rena):

A customer has an offer for 5 portrait sheets for 9.99 with additional portrait sheets priced at 19.99 each. If the customer desires to purchase 10 portrait sheets, how much would she save if she instead purchases a 10 portrait sheet package for $99.99? I got$19.95 just want to make sure thats correct....Math is NOT my forte :(

OpenStudy (anonymous):

is it a package of 5 for $9.99 all together?

OpenStudy (rena):

yes

OpenStudy (anonymous):

she would save 9.95 bcuz: 9.99 + 5(19.99) = 109.94 then subtract the 99.99, which means she saves 9.95

OpenStudy (rena):

HOw did I get $19.95???

OpenStudy (anonymous):

no idea haha

OpenStudy (rena):

me either....

OpenStudy (rena):

A digital picture frame with a suggested retail price of $500.00 is being offered on discount. The frame was initially marked down by 40%, and the customer will receive an additional 50% off the marked down price at the register. How much will the customer pay? Lets see if I got the right answer lol

OpenStudy (anonymous):

$150

OpenStudy (rena):

I had it!!!!!!

OpenStudy (anonymous):

awesome :)

OpenStudy (rena):

If a special portrait package offer consists of 2 8x10-sized portraits, 2 5x7-sized portraits, and 8 wallet-sizes portraits, what fraction of the portraits are 5x7-sized? An answer is required?? I am HORRIBLE at fractions but I keep coming up with 2/3.......

OpenStudy (anonymous):

it would be 2 out of the total of 12 items, so reduced it is 1/6

OpenStudy (rena):

what am I doing wrong??

OpenStudy (anonymous):

how do u get 2/3?.... i can help and see what ur doing wrong

OpenStudy (rena):

I'm not completing the fraction all the way....

OpenStudy (anonymous):

?

OpenStudy (rena):

Or breaking it down all the way.

OpenStudy (anonymous):

You have 2 5x7 portaits.... how many total portraits do we have?....2 + 2 + 8 = 12 so the fractions is 2/12 = 1/6.... very straight forward

OpenStudy (rena):

yea...Thank you

OpenStudy (rena):

$100 x (1-.40) x (1-.2) = I think I got it..

OpenStudy (anonymous):

yup!